# The Power of LLM-Generated Synthetic Data for Stance Detection in Online Political Discussions

### v1

@@ 3.3 Synthetic Data-driven Query By Comittee (SQBC)
The $J$ most informative samples $\mathcal{D}^{(q)}_{\text{MInf}}\subset\mathcal{D}^{(q)}$ among the unlabelled samples are the $J$ samples with the smallest scores. In the experiments we vary $J$ to study the impact of manually labelled most informative samples.
(4) Labelling the most informative samples.
This step must be performed manually, i.e., the most informative samples selected in the previous step are labelled by a human expert. Attempting to assign labels automatically using majority vote from the synthetic samples did not work well, since the most informative samples are exactly at the boundary between the classes, leading to random and thus noisy labels.
@@ 4 Experiments
We now describe the datasets, the experiments and the results. Our experiments follow the structure of the questions Q1, Q2 and Q3 posed in the introduction. Apart from presenting numerical results for these questions, we also provide further insights by visualizing the embeddings of the synthetic and real world datasets and by showing the selection behaviour of the active learning methods.
@@ 4.1 Datasets
To evaluate our methods, we use the X-Stance dataset (Vamvas and Sennrich, 2020), a multilingual stance detection dataset, which contains a total of $67,000$ annotated comments on $194$ policy-related questions, answered by Swiss election candidates. X-stance includes comments in German ($48,600$), French ($17,200$) and Italian ($1,400$). For this paper we only included the German comments, but as mentioned above, other languages can be easily included by automatic translation. The data has been extracted from smartvote, a Swiss voting advice platform. The comments are labelled either as being in favor (positive) or against (negative) the posed question, which is a typical labelling scheme for stance detection datasets. The X-Stance dataset is split in training and testing questions, i.e,. a question in the training dataset does not appear in the test dataset. Furthermore, for each question $q$ from the training data, there are several annotated comments, which form the dataset $\mathcal{D}_{\text{train}}^{(q)}$. Similarly, for each question $q$ in the test data, the set of annotated comments is written as $\mathcal{D}_{\text{test}}^{(q)}$. To refer to the whole training dataset we write $\mathcal{D}_{\text{train}}=\cup_{q\in\mathcal{Q}}\mathcal{D}^{(q)}_{\text{train}}$.
To limit computation time, we selected 10 questions from the test dataset to evaluate our method. The selection was done to best reflect the variability of the data (see Appendix D.1): we choose questions with small amount of comments, unbalanced comments and also balanced comments. Furthermore, for 5 of the questions the majority class is favor and for the other 5 the majority class is against.
@@
To test the synthetic data-augmentation and active learning based on SQBC (see Section 3.3) we generate sets of varying sizes $M=\{200,500,1000\}$ for each of the $10$ questions. The synthetic data follows the same structure as the data from the X-stance dataset, where for a specific question $q$ we have $M$ comments and $M$ labels. Each set contains $M/2$ positive labels and $M/2$ negative labels, i.e., the synthetic data is balanced.
@@ 4.2 Experimental Setup
General setup.
For all experiments, the starting model is the pre-trained BERT base [^3] https://huggingface.co/dbmdz/bert-base-german-uncased model. We adapt the model to the stance detection task by training it on the X-Stance training dataset $\mathcal{D}_{\text{train}}$ (across all questions). We call this the Baseline, and it is characteristic for BERT-based stance detection (e.g., Vamvas and Sennrich (2020)). For fine-tuning and testing we evaluate the given model separately on 10 chosen questions from the test dataset of X-Stance for all experiments. For each question $q$ we split $\mathcal{D}_{\text{test}}^{(q)}$ into a $60/40$ train/test split (repeated with 5 different seeds to get error bars) and use the train split for fine-tuning to the given question and the test split for evaluation.
Evaluated methods.
Next, we briefly describe the different evaluated methods. Table 1 shows an overview of which datasets are used for each method/variant:
Baseline: this is the default model only trained on $\mathcal{D}_{\text{train}}$, (e.g., Vamvas and Sennrich (2020)).
   True Labels: this variant assumes that all true labels for the train split of $\mathcal{D}_{\text{test}}^{(q)}$ are available
   and thus performs no active learning. This is the case $\mathcal{D}^{(q)}_{\text{MInf}}=\mathcal{D}^{(q)}_{\text{test}}$ for each $q$ in the test set.
   SQBC: we apply our active learning approach to get the most informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
CAL: we apply Contrastive Active Learning (CAL, Margatina et al. (2021)) to get the most
    informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
   Random: as an ablation, we randomly select the most informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
Note that the "+Synth" variants additionally fine-tune the model on the synthetic data $\mathcal{D}_{\text{synth}}^{(q)}$.
Evaluation.
Our main results report the average F1 score over 10 selected questions from the test dataset evaluated on the comments from the test split. The error bars represent the average standard deviation over the 10 questions for 5 runs with different seeds. More detailed results per question are shown in Appendix E.
4.3 Experiments
For each question proposed in Section 1, we designed a specific experiment to test the impact of LLM-generated synthetic data augmentation for stance detection in online political discussions:
  • Q1: The two models Baseline and True Labels are augmented with synthetic data of varying sizes.
  • Q2: We compare the 3 active learning strategies to determine the most informative samples: CAL, SQBC and Random. To study reduction in labelling effort, we define 4 different budgets of data samples ($75\%,50\%,25\%$ and $10\%$ of $\mathcal{D}^{(q)}_{\text{test}}$) to be selected for manual labelling. These variants are also compared against Baseline and True Labels.
  • Q3: We combine fine-tuning with the synthetic data (Q1) and the (manually labelled) most informative samples (Q2). We apply the same testing procedure as in Q2, where we test the performance for different labelling budgets, while also comparing to Baseline and True Labels.
4.4 Results
@@
Figure 2: Q1: Fine-tuning the model with synthetic data improves performance for increasing dataset size: Shown are the F1-Score of Baseline (left) and True Labels (right) for increasing synthetic dataset. Even if a dataset has been fully labelled, augmenting it with synthetic data proves equally as effective.
Refer to caption Fine-tuning with most informative samples selected with synthetic data
Figure 3: Q2: SQBC selects the best samples for manual labelling and improves upon the baseline model. Increasing the synthetic data size for the selection of most informative samples only improves performance marginally, which shows that uncertain samples can already be identified with a small synthetic dataset.
@@
Figure 4: Q3: Fine-tuning with most informative samples and synthetic data surpasses the model fine-tuned with all true labels: We label less samples manually, while outperforming the model with all true labels. The reason for the performance increase can be attributed on the effects from Q1 and Q2: (i) the synthetic data helps the model learn the underlying distribution. (ii) The most informative samples improve the model where the synthetic data distribution is not expressive enough.
Q1: Fine-tuning with synthetic data reaches on average the performance of the model with all true labels.
Figure 2 shows the effectiveness of fine-tuning with synthetic data by fine-tuning Baseline and True Labels with different amounts of synthetic data $M=\{200,500,1000\}$. The baseline model with $1000$ synthetic data samples almost reaches the average performance of the model with all true labels. Thus, fine-tuning with synthetic data attains similar performance to labelling all data, without having to label any real data sample. Moreover, even if a dataset has been fully labelled, augmenting it with synthetic data improves stance detection performance even further.
Q2: Most informative samples selected by SQBC are more informative than CAL and random selection and do improve upon the baseline model.
In Figure 3, we show that we can reduce labelling effort by using synthetic data to select the most informative samples. In this setting, we only use the synthetic data for active learning and do not fine-tune with the synthetic dataset. On average, SQBC is bettern than CAL and Random. We observe, that fine-tuning only with the most informative samples might not be as effective as fine-tuning with synthetic data, since Baseline+Synth (M=1000) is slightly better than SQBC. However, the advantage of using the most informative samples is that the model is grounded on real data.
Q3: Fine-tuning with the most informative samples and additionally with synthetic data surpasses the model trained with all true labels.
In Figure 4, we show the performance of fine-tuning the baseline model with the manually labelled most informative samples and at the same time with synthetic data. This combines the experiments for Q1 and Q2. We see that on average with as little as $25\%$ of the unlabelled data, we can consistently achieve better performance than the model trained with all true labels (True Labels). For $M=1000$ synthetic samples, we observe that surprisingly random sampling outperforms the other query strategies (Random+Synth). We will discuss this in the next section.
4.5 Interpreting the role of the synthetic data
The synthetic data is at the core of our work. To get a better understanding we analyse the role of synthetic data when using it for data augmentation and for the selection of the most informative samples. To study whether the synthetic data matches the real data, we visualize embeddings of the synthetic data together with the embeddings of the real world data using T-SNE projection (van der Maaten and Hinton, 2008) in Figures 5, and 6. To gain insights into the selection mechanism for different query strategies, we add the T-SNE embeddings of the selected samples in Figures 8 and 7. Again, we discuss our insights along the questions Q1, Q2 and Q3.
Q1: The synthetic data matches the distribution of the real world data.
We visualize the positive and negative synthetic data samples together with the means (cluster centers) of the real world data in Figures 5 and 6. We observe that the synthetic data extends the real world data, which we believe is the reason why fine-tuning with synthetic data is effective in online political discussions. Also, the larger the synthetic dataset size, the more the synthetic data matches the distribution of the real world data. Furthermore, the positive and negative samples are well separated, which we attribute to having pre-trained the BERT-model on $\mathcal{D}_{\text{train}}$ of the X-Stance dataset, providing prior knowledge about the stance detection task.
Q2: SQBC selects samples that are not contained in the synthetic data distribution.
We visualize the selected samples of the SQBC and compare them to the selected samples by CAL and Random in Figures 7 and 8. We see that SQBC selects the unlabelled samples that are mostly in between the two classes of the synthetic data. This is the expected behaviour since we select the samples where the classification score is ambiguous. For random selection, the range of selected samples is broad: some similar samples between the two classes like SQBC are selected, but also within class samples that are not covered by the synthetic data set. This explains why random selection works well with a large synthetic dataset, since it further extends the decision boundary of the model. Finally, CAL selects samples similar to SQBC, but mostly tends to select samples from only one class, resulting in worse performance.
Q3: The synthetic data extends the unlabelled real data, while the most informative samples improve classification by selecting samples not well covered by the synthetic data.
Combining the insights from Q1 and Q2, we see that the synthetic data effectively extends the real world data thus providing a better estimate of the underlying distribution. When adding the most informative samples from the active learning methods, we add samples that are not in the more general synthetic data distribution thus extending the decision boundary of the classifier. We argue that this is why the combination of most informative samples and synthetic samples is so effective.

### v2

@@ 3.3 Getting the Most Informative Samples: Synthetic Query By Comittee
The $J$ most informative samples $\mathcal{D}^{(q)}_{\text{MInf}}\subset\mathcal{D}^{(q)}$ among the unlabelled samples are the $J$ samples with the smallest scores. In the experiments we vary $J$ to study the impact of manually labelled most informative samples. Finally, the most informative samples are labelled by a human expert.
@@ 4.1 Datasets
We evaluate on the German dataset of the X-Stance dataset (Vamvas & Sennrich, 2020), which contains $48,600$ annotated comments on many policy-related questions ($140$ topics), answered by Swiss election candidates. We chose the German X-Stance dataset because it to our knowledge the most comprehensive stance detection dataset with a variety of topics and comments, while also being focused on online political discussions in governmental participation processes. Known english datasets such as SemEval-2016 (Mohammad et al., 2017) or P-Stance (Küçük & Can, 2020) offer less variety and are more focused on social media discussions. The comments are labelled either as being in favor (positive) or against (negative) the posed question. The dataset is split in training and testing questions, i.e,. a question in the training dataset does not appear in the test dataset. Furthermore, for each question $q$ from the training data, there are several annotated comments, which form the dataset $\mathcal{D}_{\text{train}}^{(q)}$. Analogously, for the test data we have a set of annotated comments written as $\mathcal{D}_{\text{test}}^{(q)}$. To refer to the whole training dataset we write $\mathcal{D}_{\text{train}}=\cup_{q\in\mathcal{Q}}\mathcal{D}^{(q)}_{\text{train}}$.
For our experiments, we fine-tune all stance detection models for each question separately allowing for better performance since the data distributions can vary greatly between questions (also a common scenario in online political discussions). To limit computation time, we selected 10 questions from the test dataset to evaluate our method, that best reflect the variability of the data (see Appendix E.1). We split the datasets of these questions into Test-Train and Test-Test to fine-tune the BERT model on synthetic data and to perform active learning. That is, we use the Test-Train dataset to fine-tune the model and the Test-Test dataset to evaluate the model. The number of comments in the 10 selected test questions is shown in Table 10.
Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10
  Total-Test 233 | 267 29 | 77 55 | 141 102 | 79 181 | 88 216 | 181 281 | 97 178 | 166 49 | 130 169 | 259
  Test-Train 146 | 154 19 | 44 34 | 83 68 | 40 111 | 50 130 | 108 163 | 63 100 | 106 29 | 78 98 | 158
  Test-Test 87 | 113 10 | 33 21 | 58 34 | 39 70 | 38 86 | 73 118 | 34 78 | 60 20 | 52 71 | 101
Table 1: Number of comments in our 10 selected test questions for our experiments. The numbers are split into (favor | against) labels. We have a $60-40$ (Test-Train, Test-Test) split.
@@
For synthetic data-augmentation and active learning based on SQBC (see Section 3.3) we generate synthetic datasets of varying sizes $M=\{200,500,1000\}$ for each of the $10$ questions. The synthetic data follows the same structure as the data from the X-stance dataset, where for a specific question $q$ we have $M$ comments and $M$ labels. Each set contains $M/2$ positive labels and $M/2$ negative labels, i.e., the synthetic data is balanced. We show samples of the synthetic data in Appendix C.2.
@@ 4.2 Experimental Setup
General setup. For all experiments, we start with a pre-trained BERT base model and adapt to the stance detection task by fine-tuning on the X-Stance training dataset $\mathcal{D}_{\text{train}}$ (all questions). We call this the Baseline since it is the vanilla BERT-based stance detection (e.g., Vamvas & Sennrich (2020)).
We evaluate our methods along the lines of the questions proposed in Section 1: (Q1): we analyse the effect of fine-tuning Baseline with synthetic data and compare it to the Baseline that was only fine-tuned on $D_{\text{train}}$. (Q3): we fine-tune Baseline with synthetic data and the most informative samples. We present the baselines and our methods in the following:
Baseline methods.
  • Baseline: the default BERT model fine-tuned only on $\mathcal{D}_{\text{train}}$, (e.g., Vamvas & Sennrich (2020)).
  • True Labels: we fine-tune Baseline on the true labels of $\mathcal{D}_{\text{test}}^{(q)}$.
  • Random+Synth, CAL+Synth: we use the active learning approaches to get the most informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
Our methods.
  • Baseline+Synth: we fine-tune the Baseline on the synthetic data $\mathcal{D}_{\text{synth}}^{(q)}$.
  • True Labels+Synth: we fine-tune True Labels additionally on the synthetic data $\mathcal{D}_{\text{synth}}^{(q)}$.
  • SQBC+Synth: we apply our active learning approach to get the most informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
For fine-tuning with synthetic data only (Q1), we compare the performance of our approaches to the baselines without synthetic data. For the active learning methods (Q3), we compare to the non-active learning and active learning baselines. For further experimental details we refer to Appendix D.
Analysing the synthetic data.
To analyse the synthetic data (Q2), we visualize the BERT embeddings of the synthetic data together with the embeddings of the real world data (see Figure 3 and Appendix A). We use T-SNE (van der Maaten & Hinton, 2008) to project the embeddings, i.e., the output CLS token of the BERT model to a two-dimensional space. To assess whether the synthetic data captures the overall characteristics of the real world data and shares similar labels, we plot the individual embeddings of the synthetic data together with the means of the embeddings of the real world data. Additionally, we plot their corresponding labels. Finally, we also visualize the most informative samples selected by the different active learning methods.
4.3 Results: Evaluating the effectivenes of the synthetic data
@@
Figure 2: Q1: Fine-tuning the model with synthetic data improves performance for increasing dataset size: Shown are the F1-Score of fine-tuning with Only Synthetic Data (left) and Synthetic Data + True Labels (right) for increasing synthetic dataset size. Even if a dataset has been fully labelled, augmenting it with synthetic data proves equally as effective.
Refer to captionRefer to captionRefer to captionRefer to captionSynthetic Data AligmentMost Informative SamplesSQBCRandomCALNeg. SynthPos. SynthPos. Train(Mean)Neg. Train(Mean)Selected TrainAB
Figure 3: (Q2) Analysing the synthetic data (M=1000): The synthetic data aligns well with the real data, which is crucial for improving stance detection performance and to check for potential biases introduced by the synthetic data. SQBC selects the samples that are in between the two classes, i.e, that are the most ambiguous and informative for the model.
@@
Figure 4: (Q3) Fine-tuning with synthetic data improves stance detection, while combining most informative samples and synthetic data surpasses the baseline model fine-tuned with all true labels (   above dashed line ) using less manually labelled data: The reason for the performance increase can be attributed to two phenomena: (i) the synthetic data smoothens the decision boundary of the model making it more robust to outliers. (ii) The most informative samples improve the model where the synthetic data distribution is not expressive enough.
(Q1) LLM-generated synthetic data substantially improves stance detection. We show in Figure 2, that fine-tuning with only synthetic data improves the stance detection model. For $M=1000$ the performance almost reaches the True Labels model, indicating that we can tailor a model to a certain topic without having any data for it. Furthermore, we can still improve the model consistently when new labelled data is available as seen by the True Labels+Synth models. In Section 5.1, we also analyse stance detection with zero-shot and fine-tuning approaches on Mistral-7B, but show that these are far less effective on the X-Stance dataset than our fine-tuned BERT model. In the following, we attempt to provide an understanding as to why the synthetic data is so effective.
(Q2) The synthetic data aligns well with real world data. We compare the T-SNE projected embeddings of the synthetic and real data in Figure 3(A) (more visualisations in Appendix A). The synthetic data aligns well with the real world dataset, since the means of the training data are close and in the direction of the synthetic data. We further analyse the synthetic data in Section 5.2: the synthetic data is generally of high quality which validates the notion that it can serve as a good reference distribution for the model. From a statistical learning point of view the synthetic data can be thought of smoothening the decision boundary of the model. This also explains why fine-tuning with real samples is also very effective. Furthermore, we show that it is crucial to generate synthetic data that aligns with the given topic The insight that the synthetic data provides a good prior, serves as motivation to use the synthetic data to identify ambiguous samples that are the most informative to the model. We elaborate on this in the following section.
(Q3) Synthetic data aids in finding unlabelled samples that further improve the stance detection model by extending its decision boundary. We show the results of combining the most informative samples and synthetic data in Figure 4. Combining both, we oupterform True Labels while using only $25\%$ of the labelled data. We compare the selection strategy of the methods in Figure 3(B): Due to the k-nearest neighbours objective of SQBC, the model selects samples that are in between the two classes, which proves superior to CAL and to Random for smaller synthetic data sizes. CAL performs the worst across the board: it assumes that similar embeddings that have different outputs are ambiguous, which makes it prone to outliers in the real data, e.g., when the stance detection model misclassifies a sample. Therefore, CAL often selects samples from only one class which worsens performance. Interestingly for $M=1000$, Random outperforms both active learning methods SQBC and CAL. Random selects similar samples to SQBC, but also uniformly samples from outliers from both classes, extending the decision boundary of the model. We argue this is especially effective for larger synthetic dataset sizes where the synthetic data smoothens the decision boundary and thus mitigates the high variance introduced by the most informative samples. Thus, the model remain robust while extending the decision boundary. However, as we show in Section 5.2, the real data is quite homogenous. Therefore, with severe outliers present, Random could select these and worsen performance. This would not happen with SQBC due to its k-nearest neighbour objective.
5 Ablations
We investigate different aspects of using synthetic data for online political discussions. First, we study how well LLMs perform on classifying stance on the X-Stance dataset directly. Our assumption behind this is that while LLMs are strong at generating open ended text, they seem to have more difficulties when conditioned on a specific task combined with a narrow dataset. We also study the properties of the synthetic data by calculating the per comment entropy distribution toghether with the comment length. We then compare to the real data. Furthermore, to determine whether fine-tuning on a per topic basis is sensible, we study if the observed performance improvement is related to the content or the structure of the synthetic dataset.
5.1 Using LLMs directly for X-Stance
Cruickshank & Xian Ng (2023) and Gül et al. (2024) have shown promising results using zero-shot stance detection and fine-tuning various LLMs on common stance detection datasets such as SemEval-2016 (Mohammad et al., 2017) and P-Stance (Li et al., 2021). However, both these datasets have relatively little variety in topics, with the topics addressing more popular discussion topics such as general US-Politics. For this reason, we evaluate Mistral-7B on $D_{\text{train}}$ of X-Stance, which has a larger number of niche topics. We adopt the prompt and fine-tuning scenario (fine-tuning over 4 epochs with LoRA (Hu et al., 2021)) as in Gül et al. (2024) and use our Mistral-7B model for both zero-shot stance detection and fine-tuned stance detection.
Table 3 shows that zero-shot stance detection barely reaches the performance of the pre-trained BERT baseline. Suprisingly, fine-tuning the Mistral-7B model with $\mathcal{D}_{\text{train}}$ worsened performance even further. We tried various hyperparameter settings and fine-tuned for up to 10 epochs, more than the 4 used in Gül et al. (2024). We believe the poor performance can be attributed to a few reasons: (i) The topics in the X-Stance dataset are likely not present in the training sets of the Mistral 7-B model compared to SemEval and P-Stance which contain social media comments. (ii) Due to the smaller parameter count, the model may struggle to capture both the topic and the given comment for stance prediction. This often showed with the Mistral-7B model not giving consistent classification outputs or it would often refuse to predict stance. (iii) Furthermore, the model is not tailored to give single line response. In fact, the responses were often verbose so we also accepted answers that contained the words "favor" or "against". Better prompting strategies could improve performance, however with our findings we believe that as of now using LLMs for open-ended text generation is more effective than conditioning them to give a specific output for stance detection. We also tried other similar sized open source LLMs (Llama, Openassistant, Falcon) and found that they struggled similarly in producing consistent zero-shot classification.
F1 Score (avg. 10 Questions)
Fine-tuned LLM 0.182
Zero-shot LLM 0.419
Baseline 0.693
Baseline+Synth (M=1000) 0.723
SQBC+Synth (M=1000) 0.754
Table 2: LLM-based stance detection vs BERT-based stance detection: We compare the Mistral-7B performance to the our BERT stance detection models. We see that zero-shot stance detection barely reaches the pre-trained baselines’ performance. Fine-tuning the LLM also proved difficult where the performance of the fine-tuned model worsened. Our findings for X-Stance suggest that LLMs are good at producing open-ended text, while struggling when being prompted to give a specific stance.
F1 Score (avg. 10 Questions)
M 200 500 1000
Baseline 0.693 0.693 0.693
Baseline+Synth 0.711 0.717 0.723
Baseline+Synth
(Misaligned) 0.699 0.704 0.694
Table 3: Topic alignment is crucial for the synthetic data to be effective: To determine whether improvement with synthetic data is due to the structure or content of the synthetic data, we augment the stance detection model with misaligned synthetic data. That is, the synthetic data does not align with the question given to the stance detection model. We observe the model only performs meaningfully better, when the synthetic data aligns with the posed question.
5.2 Studying the Synthetic Data
Properties of the synthetic data. To determine the quality and diversity of the synthetic data, we calculate the entropy of each comment (entropy over words) for both the synthetic and real data and compare the interquartile ranges of the corresponding entropy distributions in Figure 5. We also determine the average comment length for the interquartile ranges. The entropy reveals information about the content of both datasets, while the length acts a surrogate for structure. We observe, that the entropies between the real and synthetic data are similar, while the average comment length of the synthetic data is longer than the real data’s. Looking at the samples in Appendix C.1, we observe that the real comments have a concise (even emotional) writing style, which is common in online political discussions. The synthetic data comments are verbose and more reserved. Thus the difference in entropy could be attributed to the synthetic data containing less “emotional” comments compared to the real data. Nonetheless, we feel the synthetic samples manage to capture the content of the real data. Interestingly, the difference in length does not seem to affect the synthetic data alignment as can be seen from Figure 3. It is clear that the synthetic data is well aligned with the real data, even though the comment length between both sets of data differs. That is, the BERT model is largely invariant to the comment length and is more sensitive to the content of the comments rather than their structure.
Content vs structure: Is per topic fine-tuning necessary? To further validate our approach of fine-tuning the model per topic, we determine whether the improved performance comes from the discussion oriented structure of the data or from the content of the data. We fine-tune the model with misaligned questions and synthetic datasets, since we believe data across different topics does not necessarily share the same representation, justifying the need for per topic fine-tuning. Table 3 shows that fine-tuning with the synthetic data is only effective when the synthetic dataset is aligned with the posed question as can be seen by Baseline+Synth. Fine-tuning with synthetic data from a different topic Baseline+Synth (Misaligned) provides little improvement, delivering performance close to the Baseline. This validates the usefulness of the generated synthetic data for the stance detection model.
Statistical Summary of Synthetic Data
Total samples 10000
[…]
Entropy of Synthetic vs Real Data
Distr. Range Avg. Entropy Avg. Length
(Synthetic | Real)
Minimum 2.25 | 2.39 3 | 8
$0\%-25\%$ 2.96 | 2.78 101 | 12
$25\%-50\%$ 3.04 | 2.91 124 | 19
$50\%-75\%$ 3.06 | 2.98 134 | 26
$75\%-100\%$ 3.11 | 3.07 142 | 35
Maximum 3.27 | 3.37 282 | 53
Entropy of the Datasets
Figure 5: Synthetic data properties: (Left) The per comment entropy of the synthetic data is similar to the entropy of the real data, where the synthetic data has a higher mean entropy than the real data, while the real data has higher variance. (Right) The generated synthetic data is of high quality, with only a few outliers. Interestingly, synthetic comments are longer than real ones, making real comments denser and synthetic ones more verbose. We argue that since the projected embeddings in Figure 3 show alignment between both datasets, BERT appears to be invariant to comment length.